# Gradient Flow Sampler-based Distributionally Robust Optimization

### v1

@@ Abstract
We propose a mathematically principled PDE gradient flow framework for distributionally robust optimization (DRO). Exploiting the recent advances in the intersection of Markov Chain Monte Carlo sampling and gradient flow theory, we show that our theoretical framework can be implemented as practical algorithms for sampling from worst-case distributions and, consequently, DRO. While numerous previous works have proposed various reformulation techniques and iterative algorithms, we contribute a sound gradient flow view of the distributional optimization that can be used to construct new algorithms. As an example of applications, we solve a class of Wasserstein and Sinkhorn DRO problems using the recently-discovered Wasserstein Fisher-Rao and Stein variational gradient flows. Notably, we also show some simple reductions of our framework recover exactly previously proposed popular DRO methods, and provide new insights into their theoretical limit and optimization dynamics. Numerical studies based on stochastic gradient descent provide empirical backing for our theoretical findings.
@@ 1 Introduction
Distributionally robust optimization (DRO) (Delage and Ye,, 2010; Kuhn et al.,, 2025) is a framework that aims to enhance the robustness of the solution to optimization problems. Since the original Wasserstein distributionally robust optimization (DRO) works by (Mohajerin Esfahani and Kuhn,, 2018; Zhao and Guan,, 2018; Gao and Kleywegt,, 2016), many subsequent works have presented variations of problems with various numerical solutions. In particular, the entropy-regularized Wasserstein DRO problem can be formulated as the penalized optimization problem:
@@
where $W_{\epsilon}$ is the entropy-regularized optimal transport (OT) divergence to be defined later, $\epsilon>0$ is the entropy regularization parameter, and $\tau>0$ is a temperature parameter (Lagrange multiplier); see later discussions. This DRO problem was considered by (Wang et al.,, 2021). If we set $\epsilon=0$, we recover the penalized version of the Wasserstein DRO as considered by (Sinha et al.,, 2017). For the mathematical analysis, we will temporarily focus on the interesting choice of $W_{\epsilon}$, while other choices such as the KL divergences can also be adapted to our framework. There is a thread of works that considered the (D)RO problems via the gradient descent-ascent dynamics, e.g., (Wang and Chizat,, 2022; Yu et al.,, 2022; García Trillos and García Trillos,, 2024; Conger et al.,, 2023). For illustration, let $V:=-\ell(\theta,x)$ and $\rho_{0}=\widehat{\rho_{N}}$. The inner maximization problem can be written in the form of a proximal problem
$\min_{\rho\in{\mathcal{P}}}\;\int V\,\;\!\mathrm{d}\rho(x)+\tfrac{1}{2\tau}W^{2}_{\epsilon}(\rho,{\rho_{0}})$ (2)
@@
$\partial_{t}\rho=-\nabla\cdot\left(\nabla\rho\cdot\nabla V\right)$ (3)
@@
    The DRO problem formulation with various ambiguity sets, e.g., Wasserstein DRO (Zhao and Guan,, 2018; Mohajerin Esfahani and Kuhn,, 2018; Gao and Kleywegt,, 2016), KL-DRO (Hu and Hong,, 2013; Ben-Tal et al.,, 2013), Sinkhorn DRO (Wang et al.,, 2021), Kernel DRO (Zhu et al.,, 2021), etc.
@@
    The optimization algorithm and analysis for solving the DRO problem, e.g., WRM (Sinha et al.,, 2017), reformulation techniques (Mohajerin Esfahani and Kuhn,, 2018), SGD (Levy et al.,, 2020), etc.
In this paper, We do not invent new DRO problem formulations or ambiguity sets (item 1 above). Instead, we propose a principled mathematical framework for analysis and novel algorithms for solving DRO (item 2 above), based on the theory of gradient flows and PDEs. As a consequence of our theoretical insights, we have invented novel algorithms such as Stein variational gradient, Wasserstein-Fisher-Rao, rejection sampler-based algorithms, etc., for solving existing DRO problems with great generality and theoretical guarantees. We provide code implementations of our algorithms in the online repository: https://github.com/ZusenXu/GFS-DRO.
Furthermore, we see an elegant connection between DRO and variational methods (in the sense of calculus of variations) such as gradient flows and PDE analysis. This connection allows us to directly apply theoretical results from those domains backed by rigorous mathematical machinery such as metric space gradient flows (Ambrosio et al.,, 2005), without inventing a new theory or ad-hoc analysis techniques.
@@ 2.1 Gradient Systems and Their Gradient Flow Equations
using the inverse of the Wasserstein Riemannian metric tensor: $\mathbb{G}_{W}^{-1}(\rho):T^{*}_{\rho}{\mathcal{M}}\to T_{\rho}{\mathcal{M}},\xi\mapsto-\nabla\cdot(\rho\nabla\xi),$ where $T_{\rho}{\mathcal{M}}$ is the tangent space of $\mathcal{M}^{+}$ at $\rho$ and $T^{*}_{\rho}{\mathcal{M}}$ the cotangent space. With those ingredients, we can formally define the gradient systems that generate gradient flow equations such as the WGF equation (4). We refer to (Mielke,, 2023) for more details.
@@
2.2 Markov Chain Monte Carlo Sampler via Gradient Flow.
@@
2.3 Entropy-regularized Wasserstein DRO.
Works such as (Sinha et al.,, 2017; Wang et al.,, 2021) considered DRO problems with rather general loss functions as they are based on general-purpose continuous optimization rather than DRO reformulation techniques for special losses. A key result from (Wang et al.,, 2021) is the characterization of the worst-case distribution $\pi_{Y}$ that solves problem (2), referred to as Sinkhorn DRO therein. It has a closed-form density given by a mixture form: $\pi_{Y}=\mathbb{E}_{x\sim\widehat{\rho_{N}}}\left[\rho^{*}_{Y|X=x}\right]$, where the conditional density is
@@
for $\widetilde{V}_{x,\tau}(y):=V(y)+\tfrac{1}{2\tau}c(y,x)$, and $Z_{x}=\int\exp\left(-\tfrac{2\tau}{\epsilon}\widetilde{V}_{x,\tau}(y)\right)dy$ is a normalizing constant.
(Wang et al.,, 2021) proposed an algorithm based on solving the dual formulation of this DRO problem. They developed a specialized two-level sampling procedure to estimate the gradient of the dual objective. Different from their dual approach, this paper develops a unified gradient flow sampler framework that directly samples from the primal worst-case distribution that solves the inner problem.
@@ 3.1 Schrödinger Bridge Formulation of DRO
The optimal marginal distribution of $Y$ is given by a mixture distribution, for some normalization constant $Z_{x}$:
@@ 3.2 Gradient Flow Sampler-based DRO
Note that the formulation on the right-hand side was also observed by (Chen et al.,, 2022) in studying the proximal sampler.
@@
where we can freely choose the dissipation geometry $\mathbb{G}$ for the gradient flow. Then, our central methodology is the following perspective of sampling from conditional distributions connected to the KL-minimization problem (9). Considering solving the inner maximization problem (1) via the following two-step sampling procedure:
@@
1:Input: Initial distribution $\rho_{0}$, e.g., empirical distribution $\rho_{0}=\widehat{\rho_{N}}$, constraint set $\Theta$, $\tau,\epsilon>0$, stepsize $\eta_{t}$
@@
4:  DRO step: $\theta^{s+1}\leftarrow\text{Proj}_{\Theta}(\theta^{s}-\eta_{s}\nabla_{\theta}\ell(\theta^{s},y^{s}))$
@@
3.2.1 Wasserstein Gradient Flow Sampler.
@@
$dX_{t}=\tfrac{2\tau}{\epsilon}\nabla\widetilde{V}_{x,\tau}(X_{t})dt+dW_{t},$
@@
$X_{t+1}=X_{t}+\eta_{t}\nabla\widetilde{V}_{x,\tau}(X_{t})+\sqrt{\eta_{t}\frac{\epsilon}{\tau}}\xi_{t}$ (12)
@@
We note the step size scaling in front of the stochastic variable $\xi^{i,s}_{t}$ is different from the vanilla ULA update rule.
@@
7:   Update sample using (12): $y_{t+1}^{i,s}=y_{t}^{i,s}+\eta\nabla\widetilde{V}_{x^{s},\tau}(y_{t}^{i,s})+\sqrt{\eta\epsilon/\tau}\xi^{i,s}_{t}$
@@
$X_{t+1}=X_{t}+\eta_{t}\nabla\widetilde{V}_{x,\tau}(X_{t})$ (13)
which coincides with the inner SGD step used in the WRM algorithm by (Sinha et al.,, 2017). Henece, their WRM algorithm is a special case (with only ODE) of our GF-DRO framework in Algorithm 2.
@@
3.2.2 Wasserstein Fisher-Rao Flow Sampler.
Consider the WFR gradient system of the energy functional $F$, i.e., the triple $\left({\mathcal{P}},F,\mathsf{W\kern-1.0ptF\kern-1.0ptR}\right)$, where $\mathsf{W\kern-1.0ptF\kern-1.0ptR}$ is the Wasserstein-Fisher-Rao metric, a.k.a. the Hellinger-Kantorovich metric restricted to the probability space [^1] Note that there are many cases of misnomer in the machine learning literature: WFR should technically be defined over the space of positive measures, not the space of probability measures; see (Mielke,, 2025) for a historical account. The latter corresponds to the spherical Hellinger-Kantorovich metric. Although their gradient flow solutions can be easily related to each other via the mass scaling. See (Mielke and Zhu,, 2025) for technical details.. Specifically, we consider the HK/WFR gradient system associated with reaction-diffusion PDE:
@@
7:   $y_{t+1}^{i,s}=y_{t}^{i,s}+\eta\nabla\widetilde{V}_{x^{s},\tau}(y_{t}^{i,s})+\sqrt{\eta\epsilon/\tau}\xi^{i,s}_{t}$
8:   $w_{t+1}^{i,s}=(w_{t}^{i,s})^{1-\epsilon\eta_{w}/(2\tau)}e^{\eta_{w}\widetilde{V}_{x^{t},\tau}(y_{t}^{i,s})}$
@@
3.2.3 Stein Variational Gradient Flow Sampler.
In an attempt to simulate the (5) using deterministic particle-based methods instead of SDEs, (Liu and Wang,, 2016) proposed the Stein variational gradient descent (SVGD) algorithm, which is a implementable discrete-time algorithmic version of (5). At time step $t$, it updates the particle locations via the following gradient descent scheme.
$X^{i}_{t+1}=X^{i}_{t}-\eta\cdot\big(\frac{1}{m}\sum_{j=1}^{m}\nabla V(X^{j}_{t})k(X^{j}_{t},X^{i}_{t})+\frac{1}{m}\sum_{j=1}^{m}\nabla_{2}k(X^{j}_{t},X^{i}_{t})\big).$
@@ 4 Analysis of Gradient Flows
For notational simplicity, we abbreviate $\widetilde{V}_{x,\tau}$ as $\widetilde{V}$ in this section and, specifically, for the case of quadratic transport cost $c(y,x)=\|x-y\|^{2}$, we define the following.
@@
    Gradient dominance/PL: There exists a constant $\lambda_{\text{PL}}>0$ such that
    $\big|\nabla\widetilde{V}\big|^{2}\geq\lambda_{\text{PL}}\widetilde{V},\quad\forall x\in\mathcal{X}$ (18)
@@
    Generalized PL for functionals of measures along GF:
    $\frac{\;\!\mathrm{d}}{\;\!\mathrm{d}t}F(\rho_{t})\geq\lambda_{\text{PL}}F(\rho_{t}),$ (19)
It is immediate that
[…]
We note that the first condition in Definition 2 does not require the function $V$ to be convex. Hence, the DRO loss $\ell(\theta,x)$ does not necessarily need to be concave in variable $x$. In such cases, we can already obtain the geodesic convexity of $F$ in the Wasserstein space, termed displacement convexity.
@@
Denote the initial sample $x^{i}\sim\widehat{\rho_{N}}$ and samples running the gradient flow sampler for time $t$ are $y^{i}_{t}\sim\rho^{*}_{Y|X=x^{i}}$. Suppose that the loss function $\ell(\theta,x)$ is $L$-smooth in $x$ and $\widetilde{V}_{x,\tau}(y):=-\ell(\theta,y)+\tfrac{1}{2\tau}c(y,x)$ satisfies the gradient dominance condition in Definition 2 with $\lambda>0$. Then, in order to generate an $\epsilon$-gradient estimate, i.e.,
@@
Results such as Proposition 1 are based on the displacement convexity or PL in the Wasserstein space. The value of our general gradient flow perspective is to let us freely choose the gradient flow geometry, not confined to Wasserstein. This is not just for theoretical considerations — the estimate such as (32) depends crucially on the (PL/convexity) constant $\lambda$, which has been a major limitation of the Langevin type samplers in the literature. Note that the speed-up is due to an advantage of the Hellinger or Fisher-Rao geometry when the warm-start initialization is given. It provides a simple insight for algorithmic design: the unbalanced WFR gradient flow can be used to speed up the sampling process for DRO. This insight is later validated in the numerical studies.

### v2

@@ Abstract
We propose a mathematically principled PDE gradient flow framework for distributionally robust optimization (DRO). Exploiting the recent advances in the intersection of Markov Chain Monte Carlo sampling and gradient flow theory, we show that our theoretical framework can be implemented as practical algorithms for sampling from worst-case distributions and, consequently, DRO. While numerous previous works have proposed various reformulation techniques and iterative algorithms, we contribute a sound gradient flow view of the distributional optimization that can be used to construct new algorithms. As an example of applications, we solve a class of Wasserstein and entropy-regularized DRO problems using the recently-discovered Wasserstein Fisher-Rao and Stein variational gradient flows. Notably, we also show some simple reductions of our framework recover exactly previously proposed popular DRO methods, and provide new insights into their theoretical limit and optimization dynamics. Numerical studies based on stochastic gradient descent provide empirical backing for our theoretical findings.
@@ 1 Introduction
Distributionally robust optimization (DRO) (Delage and Ye,, 2010; Kuhn et al.,, 2025) is a framework that aims to enhance the robustness of the solution to optimization problems. Since the original Wasserstein distributionally robust optimization (DRO) works by Mohajerin Esfahani and Kuhn, (2018); Zhao and Guan, (2018); Gao and Kleywegt, (2016), many subsequent works have presented variations of problems with various numerical solutions. In particular, the entropy-regularized Wasserstein DRO problem can be formulated as the penalized optimization problem:
@@
where $W_{\epsilon}$ is the entropy-regularized optimal transport (OT) divergence to be defined later, $\epsilon>0$ is the entropy regularization parameter, and $\tau>0$ is a temperature parameter (Lagrange multiplier); see later discussions. This DRO problem was considered by Wang et al., (2021). If we set $\epsilon=0$, we recover the penalized version of the Wasserstein DRO as considered by Sinha et al., (2017). For the mathematical analysis, we will temporarily focus on the interesting choice of $W_{\epsilon}$, while other choices such as the KL divergences can also be adapted to our framework. There is a thread of works that considered the (D)RO problems via the gradient descent-ascent dynamics, e.g., Wang and Chizat, (2022); Yu et al., (2022); García Trillos and García Trillos, (2024); Conger et al., (2023). For illustration, let $V:=-\ell(\theta,x)$ and $\rho_{0}=\widehat{\rho_{N}}$. The inner maximization problem can be written in the form of a proximal problem
$\min_{\rho\in{\mathcal{P}}}\;\int V\,\;\!\mathrm{d}\rho+\tfrac{1}{2\tau}W^{2}_{\epsilon}(\rho,{\rho_{0}})$ (2)
@@
$\partial_{t}\rho=\nabla\cdot\left(\rho\nabla V\right)$ (3)
@@
    The DRO problem formulation with various ambiguity sets, e.g., Wasserstein DRO Zhao and Guan, (2018); Mohajerin Esfahani and Kuhn, (2018); Gao and Kleywegt, (2016), KL-DRO Hu and Hong, (2013); Ben-Tal et al., (2013), Sinkhorn DRO Wang et al., (2021), Kernel DRO Zhu et al., (2021), etc.
@@
    The optimization algorithm and analysis for solving the DRO problem, e.g., WRM Sinha et al., (2017), reformulation techniques Mohajerin Esfahani and Kuhn, (2018), SGD Levy et al., (2020), etc.
In this paper, We do not invent new DRO problem formulations or ambiguity sets (item 1 above). Instead, we propose a principled mathematical framework for analysis and novel algorithms for solving DRO (item 2 above), based on the theory of gradient flows and PDEs. As a consequence of our theoretical insights, we have invented novel algorithms such as Stein variational gradient, Wasserstein-Fisher-Rao gradient, rejection sampler-based algorithms, etc., for solving existing DRO problems with great generality and theoretical guarantees. We provide code implementations of our algorithms in the online repository: https://github.com/ZusenXu/GFS-DRO.
Furthermore, we see an elegant connection between DRO and variational methods (in the sense of calculus of variations) such as gradient flows and PDE analysis. This connection allows us to directly apply theoretical results from those domains backed by rigorous mathematical machinery such as metric space gradient flows Ambrosio et al., (2005), without inventing a new theory or ad-hoc analysis techniques.
@@ 2.1 Gradient Systems and Their Gradient Flow Equations
using the inverse of the Wasserstein Riemannian metric tensor: $\mathbb{G}_{W}^{-1}(\rho):T^{*}_{\rho}{\mathcal{M}}\to T_{\rho}{\mathcal{M}},\xi\mapsto-\nabla\cdot(\rho\nabla\xi),$ where $T_{\rho}{\mathcal{M}}$ is the tangent space of $\mathcal{M}^{+}$ at $\rho$ and $T^{*}_{\rho}{\mathcal{M}}$ is the cotangent space. With those ingredients, we can formally define the gradient systems that generate gradient flow equations such as the WGF equation (4). We refer to Mielke, (2023) for more details.
@@
2.2 Markov Chain Monte Carlo Sampler via Gradient Flow
@@
2.3 Entropy-regularized Wasserstein DRO
Works such as Sinha et al., (2017); Wang et al., (2021) considered DRO problems with rather general loss functions as they are based on general-purpose continuous optimization rather than DRO reformulation techniques for special losses. A key result from Wang et al., (2021) is the characterization of the worst-case distribution $\pi_{Y}$ that solves problem (2), referred to as Sinkhorn DRO therein. It has a closed-form density given by a mixture form: $\pi_{Y}=\mathbb{E}_{x\sim\widehat{\rho_{N}}}\left[\rho^{*}_{Y|X=x}\right]$, where the conditional density is
@@ 2.3 Entropy-regularized Wasserstein DRO
for $\widetilde{V}_{x,\tau}(y):=V(y)+\tfrac{1}{2\tau}c(y,x)$, and $Z_{x}=\int\exp\left(-\tfrac{2\tau}{\epsilon}\widetilde{V}_{x,\tau}(y)\right)dy$ is a normalization constant.
Wang et al., (2021) proposed an algorithm based on solving the dual formulation of this DRO problem. They developed a specialized two-level sampling procedure to estimate the gradient of the dual objective. Different from their dual approach, this paper develops a unified gradient flow sampler framework that directly samples from the primal worst-case distribution that solves the inner problem.
@@ 3.1 Schrödinger Bridge Formulation of DRO
The optimal marginal distribution of $Y$ in (8) is given by a mixture distribution, for some normalization constant $Z_{x}$:
@@ 3.2 Gradient Flow Sampler-based DRO
Note that the formulation on the right-hand side was also observed by Chen et al., (2022) in studying the proximal sampler.
@@
where we can freely choose the dissipation geometry $\mathbb{G}$ for the gradient flow. Then, our central methodology is the following perspective of sampling from conditional distributions connected to the KL-minimization problem (9). Consider solving the inner maximization problem (1) via the following two-step sampling procedure:
@@
1:Input: Initial distribution $\rho_{0}$, e.g., empirical distribution $\rho_{0}=\widehat{\rho_{N}}$, constraint set $\Theta$, $\tau,\epsilon>0$, stepsize $r_{s}$
@@
4:  DRO step: $\theta^{s+1}\leftarrow\text{Proj}_{\Theta}(\theta^{s}-r_{s}\nabla_{\theta}\ell(\theta^{s},y^{s}))$
@@
3.2.1 Wasserstein Gradient Flow Sampler
@@
$dX_{t}=-\tfrac{2\tau}{\epsilon}\nabla\widetilde{V}_{x,\tau}(X_{t})dt+dW_{t},$
@@
$X_{t+1}=X_{t}-\eta_{t}\nabla\widetilde{V}_{x,\tau}(X_{t})+\sqrt{\eta_{t}\frac{\epsilon}{\tau}}\xi_{t}$ (12)
@@
We note the stepsize scaling in front of the stochastic variable $\xi_{t}$ is different from the vanilla ULA update rule.
@@
7:   Update sample using (12): $y_{t+1}^{i,s}=y_{t}^{i,s}-\eta\nabla\widetilde{V}_{x^{s},\tau}(y_{t}^{i,s})+\sqrt{\eta\epsilon/\tau}\xi^{i,s}_{t}$
@@
$X_{t+1}=X_{t}-\eta_{t}\nabla\widetilde{V}_{x,\tau}(X_{t})$ (13)
which coincides with the inner SGD step used in the WRM algorithm by Sinha et al., (2017). Hence, their WRM algorithm is a special case (with only ODE) of our GF-DRO framework in Algorithm 2.
@@
3.2.2 Wasserstein Fisher-Rao Flow Sampler
Consider the WFR gradient system of the energy functional $F$, i.e., the triple $\left({\mathcal{P}},F,\mathsf{W\kern-1.0ptF\kern-1.0ptR}\right)$, where $\mathsf{W\kern-1.0ptF\kern-1.0ptR}$ is the Wasserstein-Fisher-Rao metric, a.k.a. the Hellinger-Kantorovich metric restricted to the probability space [^1] Note that there are many cases of misnomer in the machine learning literature: WFR should technically be defined over the space of positive measures, not the space of probability measures; see Mielke, (2025) for a historical account. The latter corresponds to the spherical Hellinger-Kantorovich metric. Although their gradient flow solutions can be easily related to each other via the mass scaling. See Mielke and Zhu, (2025) for technical details.. Specifically, we consider the HK/WFR gradient system associated with reaction-diffusion PDE:
@@ 3.2.2 Wasserstein Fisher-Rao Flow Sampler
7:   $y_{t+1}^{i,s}=y_{t}^{i,s}-\eta\nabla\widetilde{V}_{x^{s},\tau}(y_{t}^{i,s})+\sqrt{\eta\epsilon/\tau}\xi^{i,s}_{t}$
8:   $w_{t+1}^{i,s}=(w_{t}^{i,s})^{1-\epsilon\eta_{w}/(2\tau)}e^{-\eta_{w}\widetilde{V}_{x^{t},\tau}(y_{t}^{i,s})}$
@@
3.2.3 Stein Variational Gradient Flow Sampler
In an attempt to simulate the (5) using deterministic particle-based methods instead of SDEs, (Liu and Wang,, 2016) proposed the Stein variational gradient descent (SVGD) algorithm, which is a implementable discrete-time algorithmic version of (5). For a target $\pi$, at time step $t$, it updates the particle locations via the following gradient descent scheme.
$X^{i}_{t+1}=X^{i}_{t}+\eta\cdot\big(\frac{1}{m}\sum_{j=1}^{m}\nabla\log\pi(X^{j}_{t})k(X^{j}_{t},X^{i}_{t})+\frac{1}{m}\sum_{j=1}^{m}\nabla_{2}k(X^{j}_{t},X^{i}_{t})\big).$
@@ 4 Analysis of Gradient Flows
For notational simplicity, we abbreviate $\widetilde{V}_{x,\tau}$ as $\widetilde{V}$ in this section and, specifically, for the case of quadratic transport cost $c(y,x)=\|x-y\|^{2}$, we define the following. Recall the functional $F(\rho):=\frac{\epsilon}{2\tau}\operatorname{KL}\left(\rho\ \bigg|\tfrac{1}{Z_{x}}\exp\left[-\tfrac{2\tau V(y)+c(y,x)}{\epsilon}\right]\right)$ as in (11).
@@
    Gradient dominance/Polyak-Łojaciewicz (PL) inequality: There exists a constant $\lambda>0$ such that
    $\big|\nabla\widetilde{V}\big|^{2}\geq\lambda\widetilde{V},\quad\forall y\in\mathcal{X}$ (18)
@@
    Functional PL inequality over measures along GF:
    $\left\|\frac{\;\!\mathrm{d}}{\;\!\mathrm{d}t}F(\rho_{t})\right\|^{2}\geq\lambda F(\rho_{t}),$ (19)
    which is equivalent to the $\lambda$-logarithmic Sobolev inequality (LSI) for $\pi=\tfrac{1}{Z_{x}}\exp\left[-\tfrac{2\tau V(y)+c(y,x)}{\epsilon}\right]$,
    $\int\Bigg\|\nabla\log\dfrac{\rho}{\pi}\Bigg\|^{2}\;\!\mathrm{d}\rho\geq\lambda\operatorname{KL}(\rho|\pi)$ (20)
    along the solution $\rho_{t}$.
It is immediate that $1.\implies 2.$ and $1.\implies 3$. We note that the first condition in Definition 2 does not require the function $V$ itself to be convex. Hence, the DRO loss $\ell(\theta,x)$ does not necessarily need to be concave in variable $x$. In such cases, we can already obtain the geodesic convexity of $F$ in the Wasserstein space, termed displacement convexity.
@@
Denote the initial sample $x^{i}\sim\widehat{\rho_{N}}$ and samples running the Wasserstein gradient flow sampler for time $t$ are $y^{i}_{t}\sim\rho^{*}_{Y|X=x^{i}}$. For a fixed $\theta$, suppose that the loss function $\ell(\theta,x)$ is $L$-smooth in $x$ and either $\widetilde{V}_{x,\tau}(y):=-\ell(\theta,y)+\tfrac{1}{2\tau}c(y,x)$ is $\lambda$-convex with $\lambda>0$ or the functional $F$ satisfies the $\lambda$-LSI as in Definition 2. Then, in order to generate an $\epsilon$-gradient estimate, i.e.,
@@
Results such as Proposition 1 are based on the displacement convexity or PL/LSI in the Wasserstein space. The value of our general gradient flow perspective is to let us freely choose the gradient flow geometry, not confined to Wasserstein. This is not just for theoretical considerations — the estimate such as (33) depends crucially on the (PL/convexity) constant $\lambda$, which has been a major limitation of the Langevin type samplers in the literature. Note that the speed-up is due to an advantage of the Hellinger or Fisher-Rao geometry when the warm-start initialization is given. It provides a simple insight for algorithmic design: the unbalanced WFR gradient flow can be used to speed up the sampling process for DRO. This insight is later validated in the numerical studies.